# Guiding and Diversifying LLM-Based Story Generation via Answer Set Programming

### v1

@@ 1 Introduction
The clear complementarity of symbolic and neural approaches to story generation has led many researchers to try knitting these approaches together in various ways Martin et al. (2018); Ye et al. (2023); Kelly et al. (2023); Wang et al. (2024). Nevertheless, the rigidity of narrative planning has made it difficult to apply planners to open-domain story generation, even as “plan-and-write” approaches to story generation Yao et al. (2019) have grown in popularity on the LLM side.
@@ 2 Our Approach
Broadly speaking, we generate stories via a two-step process. First, we plan a large and diverse set of potential story outlines by generating the full set of solutions to a small answer set program that encodes many viable sequences of storytelling goals (Fig. 1). Second, we choose an outline at random from this set and write it into a complete textual story via LLM, subject to a high-level story premise provided as a short string of text (Fig. 2).
@@ 2.1 Plan: Outline Pre-Generation with ASP
In the plan phase, we make use of answer set programming (ASP) to generate diverse story outlines. ASP is widely used in procedural content generation research to generate artifacts ranging from story outlines to game levels Smith and Mateas (2011). Generally speaking, ASP-for-PCG approaches break down a generative answer set program into two components; the first part defines the basic elements that can be combined to form valid artifacts, and the second part defines constraints that restrict how these elements are allowed to be combined. In our case, we model story outlines as consisting of a sequence of scenes, each of which describes what happens at a certain time point in the story, and each of which fulfills some storytelling goal or narrative function related to what the story’s author is trying to convey.
@@
To maximize the diversity of generated outlines, we intentionally specify only a minimal set of constraints. Our example outline generator makes use of just fifteen total constraints, each consisting of no more than three lines of ASP code. Because story outlines are ultimately translated into complete stories by an LLM, ASP constraints do not necessarily have to prevent every problem that might arise in narrative function sequencing—they only need to prevent problems that the LLM is unable to “write around” in an acceptable way.
@@ 2.2.1 Translate Narrative Functions to Writing Instructions
We use a dictionary to translate each narrative function in the chosen outline to a detailed writing instruction for the LLM to follow.
@@ 3.1 Procedure
We evaluate our approach by generating parallel sets of stories using our (ASP-guided) approach and a naïve (unguided) baseline, then comparing the semantic homogeneity of each story set via a sentence embedding model. Intuitively, we can calculate a homogeneity score for a set of texts via the following procedure:
@@ 3.2 Results
In general, we found that our ASP-guided approach consistently generated more semantically diverse sets of stories than an unguided approach, regardless of premise and paragraph number (Fig. 3). Some premises benefited more from ASP guidance than others; for instance, stories generated in response to the “Cold Emu War” premise were much more diverse when ASP guidance was employed.
@@ 6 Conclusion
We have demonstrated the use of answer set programming (ASP) to guide LLM-based story generation toward greater diversity, using a concise answer set program (more flexible than a full-fledged narrative planning domain) to define a generator of valid story outlines and an LLM to expand these outlines into complete stories. We believe this approach represents a new reference point in the design space of neurosymbolic approaches to story generation, and a potentially compelling way to combine the open-domain story generation capabilities of LLMs with the greater content planning abilities of rules-based methods.

### v2

@@ 1 Introduction
The clear complementarity of symbolic and neural approaches to story generation has led many researchers to try knitting these approaches together in various ways Martin et al. (2018); Simon and Muise (2022); Kelly et al. (2023); Wang et al. (2024). Nevertheless, the rigidity of narrative planning has made it difficult to apply planners to open-domain story generation, even as “plan-and-write” approaches to story generation Yao et al. (2019) have grown in popularity on the LLM side.
@@ 2 Our Approach
Broadly speaking, we generate stories via a two-step process. First, we plan a large and diverse set of potential story outlines by generating the full set of solutions to a small handwritten answer set program that encodes many viable sequences of storytelling goals (Fig. 1). Second, we choose an outline at random from this set and write a complete textual story that follows this outline via LLM, conditioning generation on a high-level story premise defined by a short string of text (Fig. 2).
@@ 2.1 Plan: Outline Pre-Generation with ASP
In the plan phase, we make use of answer set programming (ASP) to generate diverse story outlines. ASP is a form of constraint solving that employs an answer set solver to generate all stable models Gelfond and Lifschitz (1988), or sets of facts, that satisfy a set of (generally hand-written) logical constraints Brewka et al. (2011).
ASP is widely used in procedural content generation research to generate artifacts ranging from story outlines to game levels Smith and Mateas (2011). Generally speaking, ASP-for-PCG approaches break down a generative answer set program into two components; the first part defines the basic elements that can be combined to form valid artifacts, and the second part defines constraints that restrict how these elements are allowed to be combined. In our case, we model story outlines as consisting of a sequence of scenes, each of which describes what happens at a certain time point in the story, and each of which fulfills some storytelling goal or narrative function related to what the story’s author is trying to convey.
@@
To maximize the diversity of generated outlines, we intentionally specify only a minimal set of constraints. Our example outline generator makes use of just fifteen total constraints, each consisting of no more than three lines of ASP code. Because story outlines are ultimately translated into complete stories by an LLM, ASP constraints do not necessarily have to prevent every problem that might arise in narrative function sequencing: they only need to prevent problems that the LLM is unable to acceptably “write around”, which can be identified through iterative testing and refinement of the full ASP-to-LLM story generation pipeline.
@@ 2.2.1 Translate Narrative Functions to Writing Instructions
We use a dictionary to map each narrative function in the chosen outline to a detailed, human-authored writing instruction for the LLM to follow.
@@ 3.1 Procedure
We evaluate our approach by generating parallel sets of ten stories each using our (ASP-guided) approach and a naïve (unguided) baseline, then comparing the semantic homogeneity of each story set via a sentence embedding model. Intuitively, we can calculate a homogeneity score for a set of texts via the following procedure:
@@ 3.2 Results
In general, we found that our ASP-guided approach consistently generated more semantically diverse sets of stories than an unguided approach, regardless of premise and paragraph number (Fig. 3).
Some premises benefited more from ASP guidance than others; for instance, stories generated in response to the “Cold Emu War” premise were much more diverse when ASP guidance was employed. Although the reason for these premise-level differences are not yet clear to us, they seem consistent across multiple different evaluation runs: for instance, story sets produced in response to the “cat pirate” premise consistently seem to benefit the least from ASP guidance in terms of diversity. Perhaps this is because, during pretraining, the LLM was exposed to more closely relevant training data for some of our evaluated story premises than for others; this could give the LLM an especially strong prior for what stories that respond to certain premises “should” look like, resulting in ASP guidance exerting less of an effect on overall story direction for these premises.
@@ 6 Conclusion
We have demonstrated the use of answer set programming (ASP) to guide LLM-based story generation toward greater diversity, using a concise answer set program (more flexible than a full-fledged narrative planning domain) to define a generator of valid story outlines and an LLM to expand these outlines into complete stories. We believe this approach represents a new reference point in the design space of neurosymbolic approaches to story generation, and a potentially compelling way to combine the open-domain story generation capabilities of LLMs with the greater content planning abilities of rules-based methods. We make our code available as a jumping-off point for future research in this direction. [^1] https://github.com/mkremins/spleenwort